5) [2 marks] Give a proof by contradiction that if n+2 is even then n is even
Answer #1
let's assume if n+2 is even, then n is odd

taken a value for n+2 which is even.
let's take n+2 = 14(an even number)
n = 14-2 = 12, which is also an even number.

but in the hypothesis, we assumed n to be an odd number.

so, by contradiction we proved that if n+2 is even, then n is an even number.